OpenStudy (anonymous):

solving a system of equations solve 4x + 3y = 13 x + 5y = 33

OpenStudy (anonymous):

eliminate y first

OpenStudy (anonymous):

1. multiply -4 to the 2nd equation -4(x+5y=33) implies -4x-20y=-132 2. add the 1st equation with the equation from the equation derive in 1 and solve for y. 4x+3y=13 + -4x-20y=-132 = -17y=-119 implies y=119/17=7 3. substitute y in any of the 2 equation and solve for x. x+5(7)=33 x=33-35=-2

OpenStudy (anonymous):

-20x-15y=-65 3x+15y=99

OpenStudy (anonymous):

thnx!

OpenStudy (anonymous):

-17x=34 x=-2

OpenStudy (anonymous):

then substitute in your original equation to get y